what is kinetic molecular theory of gases and the formula relation of boyles law, grahams law, charles law and daltons law?

Grade:6

1 Answers

Vikash
6 years ago
There was no theory in order to explain the behaviour of gases so, MAXWELL BOLTZMAN in 1857 suggested a theoratical model known as molecular theory of gases. It is also called microscopic theory of gases.
BOYLE’S   LAW: PV = 2/3 kT  [ PV IS CONSTANT]
GRAHAM;S  LAW:     [R/ R2= Mw/ Mw2]1/2
CHARLE’S  LAW :   [V/T= 2/3 . K/P]
ON’S  LAW :  P = 1/3. m1n1u1/V + 1/3. m2n2u2/ V     =     P+ P2